OpenStudy (anonymous):

When you see a traffic light turn red, you apply the brakes until you come to a stop. Suppose your initial speed was 10.7 m/s, and you come to rest in 34.7 m. How much time does this take? Assume constant deceleration

OpenStudy (anonymous):

u = 10.7 v = 0 a = ? t = ? s = 34.7 m v^2 - u^2 = 2as find a by above eq then v = u + at find t by above eq

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

uhm the answer I found is not correct. So i don't understand what you did there.

OpenStudy (anonymous):

i think you multiplied v & u by 2 in 1st eq you are supposed to square it

OpenStudy (anonymous):

no i did square it. My answer for the first equation was -1.65?

OpenStudy (anonymous):

yeah thats right is the time 6.488

OpenStudy (anonymous):

i got neg. 9.05 for the time. Can you show me how you got 6.488?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

v = u + at 0 = 10.7 -1.65t

OpenStudy (anonymous):

so t = \[-10.7\div-1.65 = \]

OpenStudy (anonymous):

that is 6.488

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

okay I see what I did wrong. Thank you !

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

fire away

OpenStudy (anonymous):

A volcano launches a lava bomb straight upward with an initial speed of 20 m/s. Taking upward to be the positive direction, find the speed and direction of the lava bomb after the following amounts of time have elapsed. a) 1.0 seconds (speed & velocity in m/s) b)3.1 seconds (speed & velocity in m/s)

OpenStudy (anonymous):

u = 20 a = g(grav) = -9.8 t = 1 , 3.1 find v using v = u + at

OpenStudy (anonymous):

thank you ! you are a physics genius !

OpenStudy (anonymous):

2*34,7*a=10,7^2 a=1,64 1,64*t=10,7 t=6,48